OpenStudy (anonymous):

What is the domain of this relation? {(-3, 0), (-2, -5), (-2, 6), (3, 7), (0, -17)}

OpenStudy (ivettef365):

domain are the values of x, you have (x,y) so what are the numbers of the domain?

OpenStudy (ivettef365):

http://www.purplemath.com/modules/fcns2.htm

OpenStudy (anonymous):

so just the x

OpenStudy (anonymous):

-3,-2,-2,3,0?

OpenStudy (ivettef365):

ok, good, but remember not to repeat numbers, and put them is order as follows Domain {-3,-2,0,3}

OpenStudy (anonymous):

thanks fo the help bruh

OpenStudy (ivettef365):

your welcome :)

OpenStudy (anonymous):

{(6, -6), (7, -7), (3, -3)}

OpenStudy (anonymous):

is that a function

OpenStudy (ivettef365):

it is a function when the x's are not repeated, therefore yes that is a function because you have domain {3,6,7}, now the other example on the top is not a function because you have x's with the same value. understand ?

OpenStudy (anonymous):

yea so this is NOT A FUNCTION {(Tom, 18), (Sue, 12), (Jorge, 28), (Sue, 28)}

OpenStudy (anonymous):

cuz sue

OpenStudy (ivettef365):

no because you have Sue twice on the x value, not a function

OpenStudy (anonymous):

yea thats wat i said

OpenStudy (ivettef365):

that is correct

OpenStudy (anonymous):

ok 1 more please

OpenStudy (ivettef365):

ok

OpenStudy (anonymous):

Find the indicated output for this function If h(x) = x - 2, find h(21)

OpenStudy (anonymous):

19?

OpenStudy (ivettef365):

just substitute the x for 21 so you have h(21) = 21 - 2 = 19 yes that is correct

OpenStudy (anonymous):

State the range of the numbers displayed in column x and column y. x y 1 5 2 7 3 9

OpenStudy (ivettef365):

the range are values of y did you see the website that I attached ?